One week, the low temperature on Tuesday was one more than four times the low on Monday, and the low on Wednesday was ten more than the low on Monday. If the sum of all three low temperatures is 35, what was the low temperatures for each of the three days.
Answer by ptaylor(2052)

Let x=low temp on Monday
Then (4x+1)=low temp on Tuesday
And (x+10)=low temp on Wednesday
Now we are told that the sum of these three temps is 35. So our equation to solve is:
x+(4x+1)+(x+10)=35 get rid of parens and collect like terms
6x+11=35 subtract 11 from both sides
6x+11-11=35-11 simplify
6x=24 divide both sides by 6
x=4 degrees-------------------------low on Monday
4x+1=16+1=17 degrees---------------------low on Tuesday
x+10=4+10=14 degrees-------------------------low on Wednesday
CK
4+17+14=35
35=35

Hope this helps-----ptaylor